Two fair dice are thrown independently. Three events A, B and C is defined as

follows :

A: Even face with first dice.

B: Even face with second dice.

C: Sum of the points on the two dice is odd.

Discuss the independence of events A, B and C.


1
Expert's answer
2021-06-01T16:12:05-0400

Probability for independent events A and B:

"P(A\land B)=P(A)\cdot P(B)"

We have:

"P(A)=1\/2,P(B)=1\/2"

"P(A\land B)=1/4"

So, Events A and B are independent.


"P(A\land C)=1/4"

"P(A)\cdot P(C)=1/2\cdot1/2=1/4"

So, Events A and C are independent. The same is for events B and C.


But:

"P(A\land B\land C)=0\ne P(A)\cdot P(B)\cdot P(C)"

So, Events A, B and C are dependent.